# From Tokens to Thoughts: How LLMs and Humans Trade Compression for Meaning

### v5

@@ B.9 Detailed AMI Scores per Model and Dataset
(Rosch, 1973c) bert-large-uncased 0.19453 0.2011 0.11336
(Rosch, 1975) bert-large-uncased 0.16547 0.27324 0.2216
(McCloskey & Glucksberg, 1978) bert-large-uncased 0.12003 0.15934 0.06306
(Rosch, 1973c) FacebookAI/roberta-large 0.1021 0.10666 0.03393
(Rosch, 1975) FacebookAI/roberta-large 0.12138 0.23938 0.14165
(McCloskey & Glucksberg, 1978) FacebookAI/roberta-large 0.06271 0.08873 0.03173
(Rosch, 1973c) google-t5/t5-large 0.16583 0.16855 0.03676
(Rosch, 1975) google-t5/t5-large -0.03799 0.04179 0.00758
(McCloskey & Glucksberg, 1978) google-t5/t5-large 0.06146 0.08825 0.0082
(Rosch, 1973c) google/gemma-2-27b 0.08523 0.09065 0.04158
(Rosch, 1975) google/gemma-2-27b 0.04276 0.10062 0.06244
(McCloskey & Glucksberg, 1978) google/gemma-2-27b 0.07814 0.10274 0.04364
(Rosch, 1973c) google/gemma-2-2b 0.04029 0.04107 0.01212
(Rosch, 1975) google/gemma-2-2b 0.04529 0.14844 0.07596
(McCloskey & Glucksberg, 1978) google/gemma-2-2b 0.09953 0.13593 0.06326
(Rosch, 1973c) google/gemma-2-9b 0.1222 0.12757 0.06053
(Rosch, 1975) google/gemma-2-9b 0.07841 0.16126 0.09617
(McCloskey & Glucksberg, 1978) google/gemma-2-9b 0.10879 0.13997 0.06439
(Rosch, 1973c) google/gemma-2b 0.04336 0.04616 0.01593
(Rosch, 1975) google/gemma-2b -0.00353 0.04483 0.01577
(McCloskey & Glucksberg, 1978) google/gemma-2b 0.03472 0.05484 0.02142
(Rosch, 1973c) google/gemma-7b 0.04459 0.04547 0.01052
(Rosch, 1975) google/gemma-7b -0.03055 0.02644 0.01506
(McCloskey & Glucksberg, 1978) google/gemma-7b 0.03338 0.05724 0.02176
(Rosch, 1973c) meta-llama/Llama-3.1-70B 0.03008 0.03528 0.01936
(Rosch, 1975) meta-llama/Llama-3.1-70B -0.07026 0.02636 0.00392
(McCloskey & Glucksberg, 1978) meta-llama/Llama-3.1-70B -0.04773 0.00972 0.00236
(Rosch, 1973c) meta-llama/Llama-3.1-8B 0.00473 0.00393 0.00023
(Rosch, 1975) meta-llama/Llama-3.1-8B -0.03928 0.05489 0.01884
(McCloskey & Glucksberg, 1978) meta-llama/Llama-3.1-8B -0.02671 0.02208 6.00E-05
(Rosch, 1973c) meta-llama/Llama-3.2-1B 0.01936 0.01567 0.00246
(Rosch, 1975) meta-llama/Llama-3.2-1B -0.01876 0.05663 0.00782
(McCloskey & Glucksberg, 1978) meta-llama/Llama-3.2-1B 0.03625 0.06798 0.01352
(Rosch, 1973c) meta-llama/Llama-3.2-3B 0.03757 0.03537 0.00876
(Rosch, 1975) meta-llama/Llama-3.2-3B 0.01893 0.09619 0.03193
(McCloskey & Glucksberg, 1978) meta-llama/Llama-3.2-3B 0.03914 0.07395 0.0202
(Rosch, 1973c) meta-llama/Meta-Llama-3-70B 0.02289 0.03133 0.01514
(Rosch, 1975) meta-llama/Meta-Llama-3-70B -0.06428 0.0185 0.00554
(McCloskey & Glucksberg, 1978) meta-llama/Meta-Llama-3-70B -0.04595 0.01068 0.00272
(Rosch, 1973c) meta-llama/Meta-Llama-3-8B 0.03512 0.02852 0.00225
(Rosch, 1975) meta-llama/Meta-Llama-3-8B -0.06011 0.03694 0.00676
(McCloskey & Glucksberg, 1978) meta-llama/Meta-Llama-3-8B -0.0355 0.0219 0.00676
(Rosch, 1973c) microsoft/deberta-large 0.03748 0.03909 0.01467
(Rosch, 1975) microsoft/deberta-large 0.16568 0.28993 0.20527
(McCloskey & Glucksberg, 1978) microsoft/deberta-large 0.03217 0.06175 0.03019
(Rosch, 1973c) microsoft/phi-1_5 0.02102 0.01786 0.0075
(Rosch, 1975) microsoft/phi-1_5 0.03989 0.13887 0.04305
(McCloskey & Glucksberg, 1978) microsoft/phi-1_5 0.00895 0.05215 0.00639
(Rosch, 1973c) microsoft/phi-1 0.0249 0.01698 0.00133
(Rosch, 1975) microsoft/phi-1 -0.03625 0.02811 0.00217
(McCloskey & Glucksberg, 1978) microsoft/phi-1 -0.01148 0.03085 0.00371
(Rosch, 1973c) microsoft/phi-2 0.03703 0.02968 0.00404
(Rosch, 1975) microsoft/phi-2 -0.03654 0.04227 0.03942
(McCloskey & Glucksberg, 1978) microsoft/phi-2 -0.00254 0.02531 0.00533
(Rosch, 1973c) microsoft/phi-4 0.03075 0.03043 0.01076
(Rosch, 1975) microsoft/phi-4 -0.06737 0.00092 -0.01361
(McCloskey & Glucksberg, 1978) microsoft/phi-4 -0.01789 0.02705 0.00066
(Rosch, 1973c) mistralai/Mistral-7B-v0.3 0.0425 0.03507 0.00357
(Rosch, 1975) mistralai/Mistral-7B-v0.3 -0.05018 0.01217 0.0177
(McCloskey & Glucksberg, 1978) mistralai/Mistral-7B-v0.3 -0.01264 0.03902 0.00931
(Rosch, 1973c) Qwen/Qwen1.5-0.5B 0.00148 -0.00225 0.00399
(Rosch, 1975) Qwen/Qwen1.5-0.5B -0.01538 0.04833 0.0095
(McCloskey & Glucksberg, 1978) Qwen/Qwen1.5-0.5B 0.02559 0.06023 0.00771
(Rosch, 1973c) Qwen/Qwen1.5-1.8B 0.03397 0.03232 0.01034
(Rosch, 1975) Qwen/Qwen1.5-1.8B -0.01129 0.05803 0.00683
(McCloskey & Glucksberg, 1978) Qwen/Qwen1.5-1.8B -0.00541 0.03614 0.00538
(Rosch, 1973c) Qwen/Qwen1.5-14B 0.0372 0.02738 0.0028
(Rosch, 1975) Qwen/Qwen1.5-14B -0.02604 0.05153 0.01211
(McCloskey & Glucksberg, 1978) Qwen/Qwen1.5-14B 0.00124 0.04136 0.00338
(Rosch, 1973c) Qwen/Qwen1.5-32B 0.02638 0.02436 0.00409
(Rosch, 1975) Qwen/Qwen1.5-32B -0.03413 0.02526 -0.00665
(McCloskey & Glucksberg, 1978) Qwen/Qwen1.5-32B -0.01991 0.02124 -0.00059
(Rosch, 1973c) Qwen/Qwen1.5-4B 0.03803 0.04058 0.01742
(Rosch, 1975) Qwen/Qwen1.5-4B -0.03309 0.03988 0.01678
(McCloskey & Glucksberg, 1978) Qwen/Qwen1.5-4B -0.03997 0.00548 -0.00028
(Rosch, 1973c) Qwen/Qwen1.5-72B 0.03697 0.02892 0.00144
(Rosch, 1975) Qwen/Qwen1.5-72B -0.06184 0.02213 0.0017
(McCloskey & Glucksberg, 1978) Qwen/Qwen1.5-72B -0.02022 0.02918 0.00297
(Rosch, 1973c) Qwen/Qwen2-0.5B 0.02266 0.01923 0.00662
(Rosch, 1975) Qwen/Qwen2-0.5B 0.0515 0.14571 0.04999
(McCloskey & Glucksberg, 1978) Qwen/Qwen2-0.5B 0.01508 0.04357 0.00643
(Rosch, 1973c) Qwen/Qwen2-1.5B 0.02956 0.02779 0.00544
(Rosch, 1975) Qwen/Qwen2-1.5B -0.03595 0.03443 -0.01099
(McCloskey & Glucksberg, 1978) Qwen/Qwen2-1.5B 0.01768 0.05407 0.01604
(Rosch, 1973c) Qwen/Qwen2-7B 0.06424 0.06439 0.02067
(Rosch, 1975) Qwen/Qwen2-7B 0.0333 0.09155 0.02832
(McCloskey & Glucksberg, 1978) Qwen/Qwen2-7B 0.05329 0.07599 0.01977
(Rosch, 1973c) Qwen/Qwen2.5-0.5B 0.03165 0.03291 0.01029
(Rosch, 1975) Qwen/Qwen2.5-0.5B -0.06534 -0.0196 -0.01165
(McCloskey & Glucksberg, 1978) Qwen/Qwen2.5-0.5B 0.0062 0.04191 0.0054
(Rosch, 1973c) Qwen/Qwen2.5-1.5B 0.04838 0.0489 0.0129
(Rosch, 1975) Qwen/Qwen2.5-1.5B 0.03785 0.113 0.02761
(McCloskey & Glucksberg, 1978) Qwen/Qwen2.5-1.5B 0.06166 0.08675 0.03162
(Rosch, 1973c) Qwen/Qwen2.5-3B 0.03882 0.0348 0.00465
(Rosch, 1975) Qwen/Qwen2.5-3B 0.03977 0.10821 0.04302
(McCloskey & Glucksberg, 1978) Qwen/Qwen2.5-3B 0.03416 0.07307 0.02959
(Rosch, 1973c) Qwen/Qwen2.5-7B 0.0529 0.05051 0.01605
(Rosch, 1975) Qwen/Qwen2.5-7B -0.00905 0.03227 0.01044
(McCloskey & Glucksberg, 1978) Qwen/Qwen2.5-7B 0.00222 0.02759 0.00551
B.10 Correlation between Human Typicality Judgments and LLM Internal Cluster Geometry
@@
B.11 Typicality and Cosine Similarity [RQ2]
@@
B.12 Theoretical Extreme Case Exploration for $\mathcal{L}$
@@ B.13 Compression Figures
Figure 17: Static Embeddings Achieve a more “Optimal” Compression-Meaning Trade-off by the $\mathcal{L}$ Measure. IB-RDT objective ($\mathcal{L}$) vs. $K$ across all datasets. Lower $\mathcal{L}$ indicates a more optimal balance between compression ($I(X;C)$) and semantic fidelity (distortion). Static embeddings consistently achieve lower $\mathcal{L}$ values than both human categories and contextual embeddings. The plots correspond to the three datasets in the following order: Rosch (1973a), Rosch (1975), McCloskey & Glucksberg (1978).
Figure 18: Contextual Embeddings Achieve Better-than-Human Compression-Meaning Trade-off by the $\mathcal{L}$ Measure. IB-RDT objective ($\mathcal{L}$) vs. $K$ across all datasets. Lower $\mathcal{L}$ indicates a more optimal balance between compression ($I(X;C)$) and semantic fidelity (distortion). Contextual embeddings outperform human categories but achieve higher $\mathcal{L}$ values than static embeddings. The plots correspond to the three datasets in the following order: Rosch (1973a), Rosch (1975), McCloskey & Glucksberg (1978).
B.14 $\mathcal{L}$ objective vs. downstream task performance

### v6

@@ Appendix I Datasets Polysemy
(rosch1973natural) bert-large-uncased 0.19453 0.2011 0.11336
(rosch1975cognitive) bert-large-uncased 0.16547 0.27324 0.2216
(mccloskey1978natural) bert-large-uncased 0.12003 0.15934 0.06306
(rosch1973natural) FacebookAI/roberta-large 0.1021 0.10666 0.03393
(rosch1975cognitive) FacebookAI/roberta-large 0.12138 0.23938 0.14165
(mccloskey1978natural) FacebookAI/roberta-large 0.06271 0.08873 0.03173
(rosch1973natural) google-t5/t5-large 0.16583 0.16855 0.03676
(rosch1975cognitive) google-t5/t5-large -0.03799 0.04179 0.00758
(mccloskey1978natural) google-t5/t5-large 0.06146 0.08825 0.0082
(rosch1973natural) google/gemma-2-27b 0.08523 0.09065 0.04158
(rosch1975cognitive) google/gemma-2-27b 0.04276 0.10062 0.06244
(mccloskey1978natural) google/gemma-2-27b 0.07814 0.10274 0.04364
(rosch1973natural) google/gemma-2-2b 0.04029 0.04107 0.01212
(rosch1975cognitive) google/gemma-2-2b 0.04529 0.14844 0.07596
(mccloskey1978natural) google/gemma-2-2b 0.09953 0.13593 0.06326
(rosch1973natural) google/gemma-2-9b 0.1222 0.12757 0.06053
(rosch1975cognitive) google/gemma-2-9b 0.07841 0.16126 0.09617
(mccloskey1978natural) google/gemma-2-9b 0.10879 0.13997 0.06439
(rosch1973natural) google/gemma-2b 0.04336 0.04616 0.01593
(rosch1975cognitive) google/gemma-2b -0.00353 0.04483 0.01577
(mccloskey1978natural) google/gemma-2b 0.03472 0.05484 0.02142
(rosch1973natural) google/gemma-7b 0.04459 0.04547 0.01052
(rosch1975cognitive) google/gemma-7b -0.03055 0.02644 0.01506
(mccloskey1978natural) google/gemma-7b 0.03338 0.05724 0.02176
(rosch1973natural) meta-llama/Llama-3.1-70B 0.03008 0.03528 0.01936
(rosch1975cognitive) meta-llama/Llama-3.1-70B -0.07026 0.02636 0.00392
(mccloskey1978natural) meta-llama/Llama-3.1-70B -0.04773 0.00972 0.00236
(rosch1973natural) meta-llama/Llama-3.1-8B 0.00473 0.00393 0.00023
(rosch1975cognitive) meta-llama/Llama-3.1-8B -0.03928 0.05489 0.01884
(mccloskey1978natural) meta-llama/Llama-3.1-8B -0.02671 0.02208 6.00E-05
(rosch1973natural) meta-llama/Llama-3.2-1B 0.01936 0.01567 0.00246
(rosch1975cognitive) meta-llama/Llama-3.2-1B -0.01876 0.05663 0.00782
(mccloskey1978natural) meta-llama/Llama-3.2-1B 0.03625 0.06798 0.01352
(rosch1973natural) meta-llama/Llama-3.2-3B 0.03757 0.03537 0.00876
(rosch1975cognitive) meta-llama/Llama-3.2-3B 0.01893 0.09619 0.03193
(mccloskey1978natural) meta-llama/Llama-3.2-3B 0.03914 0.07395 0.0202
(rosch1973natural) meta-llama/Meta-Llama-3-70B 0.02289 0.03133 0.01514
(rosch1975cognitive) meta-llama/Meta-Llama-3-70B -0.06428 0.0185 0.00554
(mccloskey1978natural) meta-llama/Meta-Llama-3-70B -0.04595 0.01068 0.00272
(rosch1973natural) meta-llama/Meta-Llama-3-8B 0.03512 0.02852 0.00225
(rosch1975cognitive) meta-llama/Meta-Llama-3-8B -0.06011 0.03694 0.00676
(mccloskey1978natural) meta-llama/Meta-Llama-3-8B -0.0355 0.0219 0.00676
(rosch1973natural) microsoft/deberta-large 0.03748 0.03909 0.01467
(rosch1975cognitive) microsoft/deberta-large 0.16568 0.28993 0.20527
(mccloskey1978natural) microsoft/deberta-large 0.03217 0.06175 0.03019
(rosch1973natural) microsoft/phi-1_5 0.02102 0.01786 0.0075
(rosch1975cognitive) microsoft/phi-1_5 0.03989 0.13887 0.04305
(mccloskey1978natural) microsoft/phi-1_5 0.00895 0.05215 0.00639
(rosch1973natural) microsoft/phi-1 0.0249 0.01698 0.00133
(rosch1975cognitive) microsoft/phi-1 -0.03625 0.02811 0.00217
(mccloskey1978natural) microsoft/phi-1 -0.01148 0.03085 0.00371
(rosch1973natural) microsoft/phi-2 0.03703 0.02968 0.00404
(rosch1975cognitive) microsoft/phi-2 -0.03654 0.04227 0.03942
(mccloskey1978natural) microsoft/phi-2 -0.00254 0.02531 0.00533
(rosch1973natural) microsoft/phi-4 0.03075 0.03043 0.01076
(rosch1975cognitive) microsoft/phi-4 -0.06737 0.00092 -0.01361
(mccloskey1978natural) microsoft/phi-4 -0.01789 0.02705 0.00066
(rosch1973natural) mistralai/Mistral-7B-v0.3 0.0425 0.03507 0.00357
(rosch1975cognitive) mistralai/Mistral-7B-v0.3 -0.05018 0.01217 0.0177
(mccloskey1978natural) mistralai/Mistral-7B-v0.3 -0.01264 0.03902 0.00931
(rosch1973natural) Qwen/Qwen1.5-0.5B 0.00148 -0.00225 0.00399
(rosch1975cognitive) Qwen/Qwen1.5-0.5B -0.01538 0.04833 0.0095
(mccloskey1978natural) Qwen/Qwen1.5-0.5B 0.02559 0.06023 0.00771
(rosch1973natural) Qwen/Qwen1.5-1.8B 0.03397 0.03232 0.01034
(rosch1975cognitive) Qwen/Qwen1.5-1.8B -0.01129 0.05803 0.00683
(mccloskey1978natural) Qwen/Qwen1.5-1.8B -0.00541 0.03614 0.00538
(rosch1973natural) Qwen/Qwen1.5-14B 0.0372 0.02738 0.0028
(rosch1975cognitive) Qwen/Qwen1.5-14B -0.02604 0.05153 0.01211
(mccloskey1978natural) Qwen/Qwen1.5-14B 0.00124 0.04136 0.00338
(rosch1973natural) Qwen/Qwen1.5-32B 0.02638 0.02436 0.00409
(rosch1975cognitive) Qwen/Qwen1.5-32B -0.03413 0.02526 -0.00665
(mccloskey1978natural) Qwen/Qwen1.5-32B -0.01991 0.02124 -0.00059
(rosch1973natural) Qwen/Qwen1.5-4B 0.03803 0.04058 0.01742
(rosch1975cognitive) Qwen/Qwen1.5-4B -0.03309 0.03988 0.01678
(mccloskey1978natural) Qwen/Qwen1.5-4B -0.03997 0.00548 -0.00028
(rosch1973natural) Qwen/Qwen1.5-72B 0.03697 0.02892 0.00144
(rosch1975cognitive) Qwen/Qwen1.5-72B -0.06184 0.02213 0.0017
(mccloskey1978natural) Qwen/Qwen1.5-72B -0.02022 0.02918 0.00297
(rosch1973natural) Qwen/Qwen2-0.5B 0.02266 0.01923 0.00662
(rosch1975cognitive) Qwen/Qwen2-0.5B 0.0515 0.14571 0.04999
(mccloskey1978natural) Qwen/Qwen2-0.5B 0.01508 0.04357 0.00643
(rosch1973natural) Qwen/Qwen2-1.5B 0.02956 0.02779 0.00544
(rosch1975cognitive) Qwen/Qwen2-1.5B -0.03595 0.03443 -0.01099
(mccloskey1978natural) Qwen/Qwen2-1.5B 0.01768 0.05407 0.01604
(rosch1973natural) Qwen/Qwen2-7B 0.06424 0.06439 0.02067
(rosch1975cognitive) Qwen/Qwen2-7B 0.0333 0.09155 0.02832
(mccloskey1978natural) Qwen/Qwen2-7B 0.05329 0.07599 0.01977
(rosch1973natural) Qwen/Qwen2.5-0.5B 0.03165 0.03291 0.01029
(rosch1975cognitive) Qwen/Qwen2.5-0.5B -0.06534 -0.0196 -0.01165
(mccloskey1978natural) Qwen/Qwen2.5-0.5B 0.0062 0.04191 0.0054
(rosch1973natural) Qwen/Qwen2.5-1.5B 0.04838 0.0489 0.0129
(rosch1975cognitive) Qwen/Qwen2.5-1.5B 0.03785 0.113 0.02761
(mccloskey1978natural) Qwen/Qwen2.5-1.5B 0.06166 0.08675 0.03162
(rosch1973natural) Qwen/Qwen2.5-3B 0.03882 0.0348 0.00465
(rosch1975cognitive) Qwen/Qwen2.5-3B 0.03977 0.10821 0.04302
(mccloskey1978natural) Qwen/Qwen2.5-3B 0.03416 0.07307 0.02959
(rosch1973natural) Qwen/Qwen2.5-7B 0.0529 0.05051 0.01605
(rosch1975cognitive) Qwen/Qwen2.5-7B -0.00905 0.03227 0.01044
(mccloskey1978natural) Qwen/Qwen2.5-7B 0.00222 0.02759 0.00551
Appendix N Correlation between Human Typicality Judgments and LLM Internal Cluster Geometry
@@
Appendix O Typicality and Cosine Similarity [RQ2]
@@
Appendix P Theoretical Extreme Case Exploration for $\mathcal{L}$
@@
Figure 27: Static Embeddings Achieve a more “Optimal” Compression-Meaning Trade-off by the $\mathcal{L}$ Measure. IB-RDT objective ($\mathcal{L}$) vs. $K$ across all datasets. Lower $\mathcal{L}$ indicates a more optimal balance between compression ($I(X;C)$) and semantic fidelity (distortion). Static embeddings consistently achieve lower $\mathcal{L}$ values than both human categories and contextual embeddings. The plots correspond to the three datasets in the following order: rosch1973internal, rosch1975cognitive, mccloskey1978natural.
Figure 28: Contextual Embeddings Achieve Better-than-Human Compression-Meaning Trade-off by the $\mathcal{L}$ Measure. IB-RDT objective ($\mathcal{L}$) vs. $K$ across all datasets. Lower $\mathcal{L}$ indicates a more optimal balance between compression ($I(X;C)$) and semantic fidelity (distortion). Contextual embeddings outperform human categories but achieve higher $\mathcal{L}$ values than static embeddings. The plots correspond to the three datasets in the following order: rosch1973internal, rosch1975cognitive, mccloskey1978natural.
Appendix R Complexity-Distortion Ratio (on the importance of $\beta$)
(a) Complexity-distortion ratio for rosch1973internal.
(b) Complexity-distortion ratio for rosch1975cognitive.
(c) Complexity-distortion ratio for mccloskey1978natural.
Figure 29: Complexity-Distortion Ratios Show That Encoder-Models are Less Sensitive to Variations in $\beta$. For each dataset, we compute the ratio between distortion (loss of human-aligned conceptual information) and complexity (representation size) across values of the rate-distortion tradeoff parameter $\beta$. Encoder models yield consistently flatter profiles, indicating that their token embeddings preserve conceptual structure even under increasing compression. Decoder models exhibit more pronounced shifts, suggesting that they redistribute representational capacity more aggressively as $\beta$ increases, leading to higher sensitivity in this tradeoff.
Figure 29 provides an additional sensitivity analysis in which we examine the ratio between the distortion and complexity components of $\mathcal{L}$ as $\beta$ varies. Across all three datasets, encoder models maintain flat profiles, indicating stable conceptual structure under compression, whereas decoder models exhibit stronger shifts, reflecting greater reallocation of representational capacity.
Appendix S $\mathcal{L}$ objective vs. downstream task performance